Activity at the Soufriere Hills Volcano during this week continued at a slightly less elevated level than during the previous week. Rockfall and associated ash clouds continues to be produced with the ash clouds being blown on the wind towards the west.
The week started with the same seismicity pattern as seen during the previous period. The small, repetitive hybrid earthquakes continued to occur at a rate of about one every two or three minutes but with reduced amplitude. After 10:00 on 26 April, the rate of occurrence of the hybrid events increased to about one per minute until approximately 20:00 on 26 April. After this time, the rate gradually decreased to about one every three to four minutes and almost completely disappeared by 29 April.
Several signals interpreted as representing small- to moderate-sized rockfalls were recorded during the whole week, with variable numbers per day. The most significant of these occurred on 26 April at 12:01, and 28 April at 07:09. These rockfalls resulted in the generation of ash clouds which were blown on the wind towards the west of the volcano, depositing small amounts of ash on Plymouth and environs. Some of these events (26 April 12:01 and 28 April 07:09) also generated small pyroclastic flows down the Tar River Valley area.
Intermittent low to moderate amplitude broadband tremor was recorded at the seismic stations closest to the volcano throughout most of the week.
A few volcano-tectonic earthquakes were located mainly at shallow a depths beneath southern Montserrat.
During the period 28 April to 01 May, several small long period earthquakes were recorded daily and located beneath the crater at shallower depths.
Several very small earthquakes have been recorded by the Gages seismic station (MGAT) during this period. The cause of these small earthquakes is not known, but appears to be localised in the Gages area. Swarms of similar events have been identified in previous records from the Gages seismic station, especially during July and August 1995.
Deformation measurements and observations
The eastern (Long Ground - Whites - Castle Peak) EDM triangle and southern (Ogarros - Chances Peak - Galways) EDM triangle were reoccupied four times and one time respectively during the week. Except for the reflector on the upper flank of Castle Peak dome for which the lines to White's Yard and Long Ground continue to show the slow shortening trend of about 1 mm per day observed since late November 1995, none of the other EDM lines show any consistent pattern of change.
Surveys of the newly established GPS network were done and the changes in line lengths are generally within the error of the technique.
The height to the top of the dome was measured at 2940 ft on 30 April, an increase of 100 ft since the last measurement on 19 April.
The Long Ground tiltmeter did not detect any major deformation events.
Correlation spectrometer surveys with a car which were started on ... April 1996 were repeated on several days during the week. The results indicated SO2 fluxes of the order of 50 to 200 tonnes per day. A number of gravity measurements, which would be used as a reference for future repeated gravity measurements, were also made using a gravimeter brought by the COSPEC team.
Visual observations of the crater area from various points on the ground around the volcano and from the helicopter were possible at most times during the week. Activity continues to be dominated by rockfalls which generate small ash clouds which are blown on the wind to the west. The area of highest rockfall activity was the northeastern, northern and eastern sectors of the dome.
Vigorous steam emission continues from various areas of the dome, especially from the southern dome.
Activity at the Soufriere Hills Volcano continued at a slightly less elevated level. Small- to moderate-sized rockfalls continue to dominate activity. Visual observations indicate that the areas of highest rockfall activity were the northern, northeastern and eastern sectors of the dome. Small long period earthquakes and intermittent low to moderate amplitude broadband tremor was recorded at seismic stations throughout the week. Ground deformation continues at a slow rate and surveys of the new GPS network indicate changes in line length which are generally within the error of the technique. COSPEC measurements have indicated SO2 fluxes of 50 to 200 tonnes per day.